The electric potential at a position located a distance of 19.2 mm from a positive point charge of 6.50×10-9C and 12.1 mm from a second point charge is 1.30 kV. Calculate the value of the second charge.
Answer #1

Solution)

potential V = 1.30*10^3 V

charge q1 = 6.5*10^-9 C

distance from charge q1 , r1 = 19.2 mm = 19.2*10^-3 m

distance from charge q2 , r2 = 12.1 mm = 12.1 *10^-3 m

Total potential V = k[q1/r1 + q2/r2]

here , constant k = 9*109 N.m2/C2

substitute values, we get

1.3*10^3 = (9*10^9)[6.5*10^-9/19.2*10^-3 + q2/12.1*10^-3]

charge q2 = -2.34*10^-9 C (Ans)

========

Good luck!:)